OpenStudy (anonymous):

simplify -4x^2-3x-(x-x^2+3)

OpenStudy (anonymous):

\[-4x^2 - 3x - (x-x^2 + 3)\]

OpenStudy (anonymous):

-3x^2-4x-3

OpenStudy (anonymous):

-3x^2-4x-3

OpenStudy (anonymous):

one way to start is to apply the "-" to the items in the parantheses

OpenStudy (anonymous):

-3x^2 -4x-3

OpenStudy (anonymous):

so you have \[-4x^2 - 3x - (x-x^2 + 3)\] = \[-4x^2 - 3x + -(x-x^2 + 3)\] = \[-4x^2 - 3x + (-x) + x^2 -3\]

OpenStudy (anonymous):

so now group like terms, x^2, x and constants

OpenStudy (anonymous):

\[-4x^2 + x^2 - 3x -x -3\]

OpenStudy (anonymous):

combine like terms

OpenStudy (anonymous):

\[-4x^2 + x^2 - 3x -x -3\] = \[(-4x^2 + x^2) + (-3x -x) -3\] = \[-3x^2 -4x - 3\]

OpenStudy (anonymous):

there you have it